OpenStudy (anonymous):

Create your own binomial expression with a radical in the second term. Part 1: Identify its conjugate and explain, in complete sentences, why it is the conjugate. (1 point) Part 2: Multiply your original binomial expression and its conjugate. What happened to the radicals and why? (1 point)

OpenStudy (anonymous):

please help me

OpenStudy (mertsj):

\[2+\sqrt{3}\]

OpenStudy (mertsj):

The conjugate is \[2-\sqrt{3}\] because if the two are multiplied together, the radical is eliminated.

OpenStudy (mertsj):

\[(2-\sqrt{3})(2+\sqrt{3})= 4+2\sqrt{3}-2\sqrt{3}-3 =1\]

OpenStudy (mertsj):

The radicals were eliminated because the outer and inner terms are opposites and the last term is the square of a square root which eliminates the radical.

jhonyy9 (jhonyy9):

- in cases like this us always formula : \[a ^{2}-b ^{2}=(a-b)(a+b)\] - and hence ,on this way will be eliminated the radical

jhonyy9 (jhonyy9):

- so again this is exactly like in cases of complex numbers when we us the conjugate for one denominator without i